Operations Report Nov 3rd

Crew 214 Operations Report 3-11-2019

SOL: 7

Name of person filing report: Andrew Wheeler

Non-nominal systems: Opportunity, Suit 10

Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not charging, Suit 5 helmet is coming apart

Generator: run

Hours run: 11h 40m

From what time last night: 1900

To what time this morning: 0740

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before generator is run at night): 94%

Diesel Reading – 65%

Station Propane Reading – 83%

Ethanol Free Gasoline: N/A

Water (loft tank): 47 gallons

Water Meter: 1455453 units

Water (static tank): 246 gallons

Static to Loft Pump used – yes

Water in Green Hab: 225 gallons

Water in Science Dome: 0 gallons

Toilet tank emptied: no

Deimos rover used: not at hab

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: –

Beginning charge: –

Ending charge: –

Currently charging: –

Spirit rover used: no

Hours: (before EVA): 107.0

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 100%

Currently charging: yes

Opportunity rover used: no

Hours: 65.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Curiosity rover used: no

Hours: 113.4

Beginning charge: (Before EVA) : 100%

Ending charge: (On return from EVA, before charging): 100%

Currently charging: yes

Notes on rovers: no

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: N/A

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were used today: 0

Notes on ATVs: Assigned to the director

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: Functional

Summary of suits and radios: 4 radios not functional, Suit 10 not charging, need 6 extra batteries, Suit 5 helmet is coming apart

Summary of Hab operations: All systems nominal

Summary of GreenHab operations: Watering and fertilization of the soil commenced and 14 plants were planted.

Summary of ScienceDome operations: All systems nominal.

Summary of RAM operations: Not used

Summary of any observatory issues: N/A

Summary of health and safety issues: All crew healthy

Questions, concerns and requests to Mission Support: Still need batteries for 6 radios, suit 5 helmet needs minor fix.

Copyright © The Mars Society. All rights reserved. | Main Site